# JWST PRIMER: A lack of outshining in four normal galaxies from the ALMA-CRISTAL Survey

### v1

@@ 2.3 Sample
In this study we focus on a sample of spectroscopically confirmed, massive Lyman-break galaxies at ā that are both in the CRISTAL sample and have deep multi-band NIRCam imaging from the PRIMER survey. The sources were initially part of a large -band selected sample in COSMOS, of which 118 were observed with ALMA as part of the ALPINE large programĀ (Le FĆØvre etĀ al., 2020). They were originally spectroscopically confirmed via the detection of the Lyman- line111The original IDs as part of this program, and carried into the ALMA ALPINE follow-up are DEIMOS_COSMOS_630594 (C11), vuds_cosmos_5100994794 (C13), vuds_cosmos_5101244930 (C15), DEIMOS_COSMOS_742174 (C17), and subsequently detected in the [CII]Ā cooling line in ALPINE to provide a systemic redshift (see TableĀ 1). Several CRISTAL galaxies haveĀ JWST imaging or spectroscopic follow-up, however only four benefit from the full eight NIRCam filters from PRIMER. Postage stamp images of these galaxies are shown in Fig.Ā 1.
@@
As detailed inĀ Le FĆØvre etĀ al. (2020) andĀ Faisst etĀ al. (2020), the ALPINE sample includes galaxies at selected using different techniques, with spectroscopic follow-up including sources with both rest-frame UV emission and absorption features. Using the [CII]Ā detections, the galaxies were given morpho-kinematic classifications as an indication of the potential formation mechanism, based on the galaxy morphology and observed kinematic structure. We present the rest-frame equivalent-widths () from the ALPINE catalogue in TableĀ 1. Cassata etĀ al. (2020) note that the Lyman- emission in CRISTAL-11 and CRISTAL-15 may be offset from the optical and [CII]Ā peak by arcsec (which is approximately kpc at this redshift).
@@
CRISTAL-11 and CRISTAL-13 were initially identified using the Lyman-break technique. They both have an assigned morpho-kinematic class of āextended dispersion-dominatedā, which corresponds to a resolved/extended [CII]Ā emission line that did not show a velocity gradient in the ALPINE [CII]Ā data. CRISTAL-13 also has a classification of āUNCā inĀ Jones etĀ al. (2021), indicating that the merger or rotating disk classications cannot be robustly assigned. In contrast, CRISTAL-15 and CRISTAL-17 were both initially selected as narrow-band detected sources, meaning that they had significant Lyman- emission to cause an excess in the narrow-band images (SuprimeCam NB711 at and NB814 at for the two sources respectively). The morpho-kinematic class of CRISTAL-15 is a āpair-mergerā, indicating that the source shows evidence for an interaction, typically showing a disturbed velocity map or separate components in the [CII]Ā or auxiliary data. As stated inĀ Cassata etĀ al. (2020), this source has multiple components in the optical data and hence was classified as a merger. Finally, CRISTAL-17 has not been allocated a morpho-kinematic class from these previous works due to a weak [CII]Ā emission line, howeverĀ Cassata etĀ al. (2020) identify that the [CII]Ā emission is offset by 0.7 arcsec from the peak of the Lyman- and optical emission.
@@ 3 Methods
Here we present the key steps in our analysis of the high spatial resolutionĀ JWST NIRCam images from PRIMER. Specifically, we carry out a pixel-by-pixel SED fitting analysis and estimate the rest-frame UV and optical sizes from a parametric and non-parametric fitting approach, which we then can compare to the dust continuum sizes obtained inĀ Mitsuhashi etĀ al. (2023).
@@ 3.1 Pixel fitting criteria
We created an image mask for each source that identified pixels in which we are confident that we can determine robust physical parameters from our SED fitting analysis. The mask was created using an inverse-variance weighted (IVW) stack, combining all the available bands. Pixels were selected if they reached a SNR of 20 in this band-combined stack. This criteria results in 100ā450 individual pixels being fit per galaxy, with the major features retained for each source across the wide wavelength range covered. The fitting of these pixels results in derived physical parameters that we call the resolved quantities following the terminology ofĀ GimĆ©nez-Arteaga etĀ al. (2023). Integrated photometry was obtained for each galaxy by summing the fluxes of each individual pixel used in the resolved fits to produce a single flux value for each photometric band. This was done (as opposed to aperture photometry) to allow for a direct comparison of the inferred parameters in a resolved versus integrated analysis using the same pixels. Both the resolved and integrated fits miss a proportion of the total galaxy light due to the conservative mask we use, which corresponds to a recovery fraction of 0.72, 0.69, 0.78 and 0.53 of the total light for CRISTAL-11, -13, -15 and -17 respectively. The total flux here was determined in each band by summing the flux in the cutout image at significance, taking an average across the NIRCam bands available (no significant trends with wavelength were seen, as supported by the comparable rest-frame UV and optical sizes we determine). This method for determining the total flux gives consistent results to using a fixed large aperture. We correct both the resolved parameters and the integrated photometry by these factors. The enclosed flux of the mask used for CRISTAL-17 contains a significantly lower fraction of the total galaxy light than the other three sources: CRISTAL-17 is the faintest galaxy in our sample and therefore a lower proportion of the pixels pass our conservative masking criterion. As we show in Section 4.2.2, CRISTAL-17 also shows a significantly lower flux in the NIRCam data compared to the previous ground-based andĀ Spitzer/IRAC photometry. We identify a lower redshift companion galaxy to the North of CRISTAL-17 that appears to have contaminated the measured flux for the central source in the previous COSMOS2015 and COSMOS2020 catalogues (most strikingly by a factor of in the Ā band), leading to the observed lower fluxes in the high resolution NIRCam data.
@@ 3.2 SED fitting analysis
We allow the age to vary from 1 Myr to 1 Gyr, and we fix the metallicity at 0.5 . The dust attenuation law is assumed to beĀ Calzetti etĀ al. (2000) and we allow the magnitude of attenuation to vary in the range 0 < < 2. We include nebular emission in our model using the BAGPIPES default parametrisation and allowing the ionisation parameter to vary in the range . Nebular and stellar light is assumed to be subject to the same attenuation. We set a floor to the error of 5 percent as is standard in SED fitting analyses. We fix the redshifts of the four objects to the previously determined systemic spectroscopic redshifts from the ALPINE programĀ (Le FĆØvre etĀ al., 2020). Applying a resolved methodology like that performed here allows us to observe regions of these galaxies that would otherwise be blended and fitted as a single source (with a single value of e.g. , age) as we discuss in detail below. We homogenised the PSF of the images to match that of coarsest spatial resolution data, which provided a smoothing of the derived results over the scale of the F444W PSF size, which is arcsec or five pixels. Therefore neighboring pixels are not independent in the fitting analysis on this scale.
The parameters inferred for each pixel fit allow us to build maps of each parameter matching the spatial resolution of the F444W image, where we display the 50th percentile from the posterior distribution derived in BAGPIPES. In addition to parameter estimates produced from the SED fitting, we also inferred other parameters from the best-fitting SED model in the integrated and resolved cases. By fitting a power-law slope in the wavelength range of āĆ Ā (Calzetti etĀ al., 1994) we extracted an estimate of the rest-frame UV slope (, where ). Rest-frame equivalent widths of H and H + [OIII]Ā were also derived from the best-fitting SED using the indices fitting function in BAGPIPES. We also present the empirical colour maps across the rest-frame UV and Balmer break region as obtained by subtracting the relevant broad-band images. For CRISTAL-11, CRISTAL-13 and CRISTAL-15 at the F090WF150W colour is a proxy for the rest-UV slope and the Balmer break is bracketed by F150WF277W, while for CRISTAL-17 at the appropriate colours are F090WF200W and F200WF356W. We note here that the rest-frame optical emission lines of H lies within the F356W band for CRISTAL-11, -13 and -15 and in F444W (not F410M) for CRISTAL-17. The H + [OIII]Ā complex lies within F277W for the three sources, and in F356W for CRISTAL-17 at .
@@ 4 Results
All four galaxies are clearly spatially resolved in the PRIMERĀ JWST images, extending over several hundred NIRCam pixels at significance in the IVW stack, as shown in Fig.Ā 1. For all of the galaxies in the sample we are able to identify at least two distinct components at the resolution ofĀ JWST. This is in contrast to the previousĀ HST imaging where the sources typically appears as a single extended object in the rest-frame UV to optical wavelengths. In this section we first present a detailed exploration of the results of the resolved SED study, followed by a comparison of the derived resolved properties to that from the integrated photometry.
@@ 4.1 Resolved SED fitting
In Fig.Ā 2 and Fig.Ā 3 we present the pixel-by-pixel fitting results. These show the BAGPIPES derived dust-corrected SFR, stellar mass, specific star-formation rate (sSFR = SFR/), mass-weighted age, and , in addition to the rest-frame UV slope and rest-frame equivalent width of the estimated H and H + [OIII]Ā emission lines. We also show empirical colour maps that probe the rest-UV slope and Balmer break to present a model independent view of the galaxy colours in these spectral regions. Reassuringly, the process we used to decide whether to include pixels in the fitting analysis provides maps that have a consistent shape to that observed in the original images (Fig.Ā 1). We include contours from the F090W images prior to PSF homogenisation as an overlay to the empirical colour maps. We also show dust continuum contours from the ALMA-CRISTAL program in the image where there was significant (>3) signal, in the case of CRISTAL-11 and CRISTAL-13. Consistent results are also found when assuming a delayed- SFH, as presented in Fig.Ā 9 and Fig.Ā 10. In general, for all of the sources we find the empirical colour map that probed the rest-frame UV corresponds well to the SED fitting derived slope , as expected. Similarly, the colour probing the Balmer break (shown by the bottom left panels in Fig.Ā 2 and Fig.Ā 3) well matches the age and stellar mass maps, giving us further reassurance that the model fitting approach is robust. Furthermore, younger populations tend to be the bluest in the rest-frame UV slope , which overlaps with the regions of the highest sSFR and rest-frame optical emission line strength as expected. We visually inspected theĀ HST/ACS F606W image obtained from the CANDELS surveyĀ (Grogin etĀ al., 2011; Koekemoer etĀ al., 2011) for all sources to verify that the different components are drop-out sources, and hence likely at the same redshift. We now discuss in detail the four sources in turn.
@@ 4.1.1 CRISTAL-11
The galaxy CRISTAL-11 (DEIMOS_COSMOS_630594, at ), appears as a single elongated source in the previousĀ HST optical and NIR data as shown in Fig.Ā 1. In NIRCam, we find that the source is formed of two distinct components, which we will call North and South in subsequent discussion. First considering the colour-colour maps for this source, we see that the rest-frame UV colour is fairly uniform and blue (F090W-F150W ) over the two components, while the colour bracketing the Balmer break (F150WF277W) shows a strong gradient of , with significantly redder colours in the North region. The derived -slope map, in agreement with the empirical colour map of F090W-F150W, shows a relatively uniform blue value over the source, with the North component being slightly redder than the South ( and in the North and South respectively). These observations suggest that the North region of CRISTAL-11 harbours potentially older stellar populations, or young stars with additional dust reddening.
Using the derived physical properties from the SED fitting analysis we can disentangle these effects and determine the role of age and dust in the observed colours of the source. The North region, as expected from the redder colours across the Balmer break, contains the peak of the stellar mass, showing a higher dust attenuation () and an older age () in comparison to the South region. Instead, the South component consists of the peak of the star formation in the galaxy, showing evidence for a young () stellar population that has very little dust attenuation. This region corresponds to the peak in SFR and sSFR. If we separate the source by the two regions, we find that the younger South component contains only around a third of the total mass, despite being the brighter component in theĀ HST data. We find the inferred HĀ and H + [OIII]Ā emission line strengths are relatively weak and the value is affected by the flux errors in the F277W and F356W bands, leading to values that are not strongly correlated with position in the galaxy. With the newĀ JWST NIRCam data extending to we can extract the centre of mass of the galaxy. We find that this is offset by approximately 0.5 arcsec (3.5 kpc) from the rest-frame UV peak. The centre of mass of the system (in the North component) is also the location of the dust continuum detected by ALMA/CRISTAL (discussed further in SectionĀ 5.4).
@@ 4.1.2 CRISTAL-13
CRISTAL-13 (vuds_cosmos_5100994794, at , shown in Fig.Ā 2, has the largest physical extent of the four CRISTAL galaxies studied in this paper. It is formed of a compact North East component, and a diffuse and clumpy tail to the West, which extends over 1.5 arcsec or 10 kpc. In the Western tail (WT), we identify at least 5 separate clumps in the higher spatial resolution F090W to F200W bands shown in Fig.Ā 1. When the data is convolved to the resolution of the F444W band for the SED fitting analysis, these regions are less easily distinguished, however we see structure in the derived parameter maps over these scales (e.g. see the contours shown in the F090W-F150W map in Fig.Ā 2). The clumps within the Western tail show correlated values of high sSFR, young age, and blue rest-frame UV slope, which are also correlated with the F090WāF150W colour as expected. We see several peaks in the derived H and H + [OIII]Ā emission in the Western tail, further corroborating that these are regions of high sSFR with young ( Myr) ages.
The NE component and the Western tail are both very blue () and embedded within regions of redder rest-frame UV colour. We verified that this halo of redder colours was present in the images directly, and not an artefact of the fitting process. The NE component is compact and shows a colour gradient such that the galaxy is reddest between the NE component and the blue tail. This NE regions is characterised by a relatively complex age and Ā structure, which is due to degeneracies between these properties in the fitting. The redder region of the NE component, which is also picked up in the Balmer (F150WāF277W) colour maps, corresponds to the centre of mass of the system, which is subtly offset from the peak in SFR. The peak in mass is however co-spatial with the peak in . Despite containing a quarter of the pixels, the wider NE component (that includes the SFR and stellar mass peaks) contains half of the total mass. CRISTAL-13 contains a significant ALMA detection of the dust continuum (8), which spatially coincides with the strongest inferred .

### v2

@@ 2.3 Sample
In this study we focus on a sample of spectroscopically confirmed, massive Lyman-break galaxies at – that are both in the CRISTAL sample and have deep multi-band NIRCam imaging from the PRIMER survey. The sources were initially part of a large -band selected sample in COSMOS, of which 118 were observed with ALMA as part of the ALPINE large program (Le Fèvre et al., 2020). They were originally spectroscopically confirmed via the detection of the Lyman- line111The original IDs as part of this program, and carried into the ALMA ALPINE follow-up are DEIMOS_COSMOS_630594 (C11), vuds_cosmos_5100994794 (C13), vuds_cosmos_5101244930 (C15), DEIMOS_COSMOS_742174 (C17), and subsequently detected in the [CII] cooling line in ALPINE to provide a systemic redshift (see Table 1). Several CRISTAL galaxies have JWST imaging or spectroscopic follow-up, however only four benefit from the full eight NIRCam filters from PRIMER. Postage stamp images of these galaxies are shown in Fig. 1.
@@
As detailed in Le Fèvre et al. (2020) and Faisst et al. (2020), the ALPINE sample includes galaxies at selected using different techniques, with spectroscopic follow-up including sources with both rest-frame UV emission and absorption features. Using the [CII] detections, the galaxies were given morpho-kinematic classifications as an indication of the potential formation mechanism, based on the galaxy morphology and observed kinematic structure. We present the rest-frame equivalent-widths () from the ALPINE catalogue in Table 1. Cassata et al. (2020) note that the Lyman- emission in CRISTAL-11 and CRISTAL-15 may be offset from the optical and [CII] peak by arcsec (which is approximately kpc at this redshift).
@@
CRISTAL-11 and CRISTAL-13 were initially identified using the Lyman-break technique. They both have an assigned morpho-kinematic class of ‘extended dispersion-dominated’, which corresponds to a resolved/extended [CII] emission line that did not show a velocity gradient in the ALPINE [CII] data. CRISTAL-13 also has a classification of ‘UNC’ in Jones et al. (2021), indicating that the merger or rotating disk classications cannot be robustly assigned. In contrast, CRISTAL-15 and CRISTAL-17 were both initially selected as narrow-band detected sources, meaning that they had significant Lyman- emission to cause an excess in the narrow-band images (SuprimeCam NB711 at and NB814 at for the two sources respectively). The morpho-kinematic class of CRISTAL-15 is a ‘pair-merger’, indicating that the source shows evidence for an interaction, typically showing a disturbed velocity map or separate components in the [CII] or auxiliary data. As stated in Cassata et al. (2020), this source has multiple components in the optical data and hence was classified as a merger. Finally, CRISTAL-17 has not been allocated a morpho-kinematic class from these previous works due to a weak [CII] emission line, however Cassata et al. (2020) identify that the [CII] emission is offset by 0.7 arcsec from the peak of the Lyman- and optical emission.
As presented in Herrera-Camus et al. 2025, in the ALMA-CRISTAL data all four galaxies have [CII] detections with a measured signal-to-noise of 3.7–12. For CRISTAL-11 and -13, the [CII] peaks at the position of the dust continuum emission, but there is additional [CII] flux extending beyond the rest-frame UV/optical emission. CRISTAL-15 has [CII] emission centered on the source, with the weak emission in CRISTAL-17 being offset to the North. From kinematic analysis presented in Lee et al. in prep, CRISTAL-11 and CRISTAL-15 show evidence for being a rotating disk.
@@ 3 Methods
Here we present the key steps in our analysis of the high spatial resolution JWST NIRCam images from PRIMER. Specifically, we carry out a pixel-by-pixel SED fitting analysis and estimate the rest-frame UV and optical sizes from a parametric and non-parametric fitting approach, which we then can compare to the dust continuum sizes obtained in Mitsuhashi et al. (2023).
@@ 3.1 Pixel fitting criteria
We created an image mask for each source that identified pixels in which we are confident that we can determine robust physical parameters from our SED fitting analysis. The mask was created using an inverse-variance weighted (IVW) stack, combining all the available bands. Pixels were selected if they reached a SNR of 20 in this band-combined stack. This criteria results in 100–450 individual pixels being fit per galaxy, with the major features retained for each source across the wide wavelength range covered. The fitting of these pixels results in derived physical parameters that we call the resolved quantities following the terminology of Giménez-Arteaga et al. (2023). Integrated photometry was obtained for each galaxy by summing the fluxes of each individual pixel used in the resolved fits to produce a single flux value for each photometric band. This was done (as opposed to aperture photometry) to allow for a direct comparison of the inferred parameters in a resolved versus integrated analysis using the same pixels. Both the resolved and integrated fits miss a proportion of the total galaxy light due to the conservative mask we use, which corresponds to a recovery fraction of 0.72, 0.69, 0.78 and 0.53 of the total light for CRISTAL-11, -13, -15 and -17 respectively. The total flux here was determined in each band by summing the flux in the cutout image at significance, taking an average across the NIRCam bands available (no significant trends with wavelength were seen, as supported by the comparable rest-frame UV and optical sizes we determine). This method for determining the total flux gives consistent results to using a fixed large aperture. We correct both the resolved parameters and the integrated photometry by these factors. The enclosed flux of the mask used for CRISTAL-17 contains a significantly lower fraction of the total galaxy light than the other three sources: CRISTAL-17 is the faintest galaxy in our sample and therefore a lower proportion of the pixels pass our conservative masking criterion. As we show in Section 4.2.2, CRISTAL-17 also shows a significantly lower flux in the NIRCam data compared to the previous ground-based and Spitzer/IRAC photometry. We identify a lower redshift companion galaxy to the North of CRISTAL-17 that appears to have contaminated the measured flux for the central source in the previous COSMOS2015 and COSMOS2020 catalogues (most strikingly by a factor of in the band), leading to the observed lower fluxes in the high resolution NIRCam data.
@@ 3.2 SED fitting analysis
We allow the age to vary from 1 Myr to 1 Gyr, and we fix the metallicity at 0.5 . The dust attenuation law is assumed to be Calzetti et al. (2000) and we allow the magnitude of attenuation to vary in the range 0 < < 2. We include nebular emission in our model using the BAGPIPES default parametrisation and allowing the ionisation parameter to vary in the range . Nebular and stellar light is assumed to be subject to the same attenuation. We set a floor to the error of 5 percent as is standard in SED fitting analyses. We fix the redshifts of the four objects to the previously determined systemic spectroscopic redshifts from the ALPINE program (Le Fèvre et al., 2020). Applying a resolved methodology like that performed here allows us to observe regions of these galaxies that would otherwise be blended and fitted as a single source (with a single value of e.g. , age) as we discuss in detail below. We homogenised the PSF of the images to match that of coarsest spatial resolution data, which provided a smoothing of the derived results over the scale of the F444W PSF size, which is arcsec or five pixels. Therefore neighboring pixels are not independent in the fitting analysis on this scale.
The parameters inferred for each pixel fit allow us to build maps of each parameter matching the spatial resolution of the F444W image, where we display the 50th percentile from the posterior distribution derived in BAGPIPES. In addition to parameter estimates produced from the SED fitting, we also inferred other parameters from the best-fitting SED model in the integrated and resolved cases. By fitting a power-law slope in the wavelength range of –Å (Calzetti et al., 1994) we extracted an estimate of the rest-frame UV slope (, where ). Rest-frame equivalent widths of H and H + [OIII] were also derived from the best-fitting SED using the indices fitting function in BAGPIPES. We also present the empirical colour maps across the rest-frame UV and Balmer break region as obtained by subtracting the relevant broad-band images. For CRISTAL-11, CRISTAL-13 and CRISTAL-15 at the F090WF150W colour is a proxy for the rest-UV slope and the Balmer break is bracketed by F150WF277W, while for CRISTAL-17 at the appropriate colours are F090WF200W and F200WF356W. We note here that the rest-frame optical emission lines of H lies within the F356W band for CRISTAL-11, -13 and -15 and in F444W (not F410M) for CRISTAL-17. The H + [OIII] complex lies within F277W for the three sources, and in F356W for CRISTAL-17 at .
@@ 4 Results
All four galaxies are clearly spatially resolved in the PRIMER JWST images, extending over several hundred NIRCam pixels at significance in the IVW stack, as shown in Fig. 1. For all of the galaxies in the sample we are able to identify at least two distinct components at the resolution of JWST. This is in contrast to the previous HST imaging where the sources typically appears as a single extended object in the rest-frame UV to optical wavelengths. In this section we first present a detailed exploration of the results of the resolved SED study, followed by a comparison of the derived resolved properties to that from the integrated photometry.
@@ 4.1 Resolved SED fitting
In Fig. 2 and Fig. 3 we present the pixel-by-pixel fitting results. These show the BAGPIPES derived dust-corrected SFR, stellar mass, specific star-formation rate (sSFR = SFR/), mass-weighted age, and , in addition to the rest-frame UV slope and rest-frame equivalent width of the estimated H and H + [OIII] emission lines. We also show empirical colour maps that probe the rest-UV slope and Balmer break to present a model independent view of the galaxy colours in these spectral regions. Reassuringly, the process we used to decide whether to include pixels in the fitting analysis provides maps that have a consistent shape to that observed in the original images (Fig. 1). We include contours from the F090W images prior to PSF homogenisation as an overlay to the empirical colour maps. We also show dust continuum contours from the ALMA-CRISTAL program in the image where there was significant (>3) signal, in the case of CRISTAL-11 and CRISTAL-13. Consistent results are also found when assuming a delayed- SFH, as presented in Fig. 9 and Fig. 10. In general, for all of the sources we find the empirical colour map that probed the rest-frame UV corresponds well to the SED fitting derived slope , as expected. Similarly, the colour probing the Balmer break (shown by the bottom left panels in Fig. 2 and Fig. 3) well matches the age and stellar mass maps, giving us further reassurance that the model fitting approach is robust. Furthermore, younger populations tend to be the bluest in the rest-frame UV slope , which overlaps with the regions of the highest sSFR and rest-frame optical emission line strength as expected. We visually inspected the HST/ACS F606W image obtained from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011) for all sources to verify that the different components are drop-out sources, and hence likely at the same redshift. We now discuss in detail the four sources in turn.
@@ 4.1.1 CRISTAL-11
The galaxy CRISTAL-11 (DEIMOS_COSMOS_630594, at ), appears as a single elongated source in the previous HST optical and NIR data as shown in Fig. 1. In NIRCam, we find that the source is formed of two distinct components, which we will call North and South in subsequent discussion. First considering the colour-colour maps for this source, we see that the rest-frame UV colour is fairly uniform and blue (F090W-F150W ) over the two components, while the colour bracketing the Balmer break (F150WF277W) shows a strong gradient of , with significantly redder colours in the North region. The derived -slope map, in agreement with the empirical colour map of F090W-F150W, shows a relatively uniform blue value over the source, with the North component being slightly redder than the South ( and in the North and South respectively). These observations suggest that the North region of CRISTAL-11 harbours potentially older stellar populations, or young stars with additional dust reddening.
Using the derived physical properties from the SED fitting analysis we can disentangle these effects and determine the role of age and dust in the observed colours of the source. The North region, as expected from the redder colours across the Balmer break, contains the peak of the stellar mass, showing a higher dust attenuation () and an older age () in comparison to the South region. Instead, the South component consists of the peak of the star formation in the galaxy, showing evidence for a young () stellar population that has very little dust attenuation. This region corresponds to the peak in SFR and sSFR. If we separate the source by the two regions, we find that the younger South component contains only around a third of the total mass, despite being the brighter component in the HST data. We find the inferred H and H + [OIII] emission line strengths are relatively weak and the value is affected by the flux errors in the F277W and F356W bands, leading to values that are not strongly correlated with position in the galaxy. With the new JWST NIRCam data extending to we can extract the centre of mass of the galaxy. We find that this is offset by approximately 0.5 arcsec (3.5 kpc) from the rest-frame UV peak. The centre of mass of the system (in the North component) is also the location of the dust continuum detected by ALMA/CRISTAL (discussed further in Section 5.4).
@@ 4.1.2 CRISTAL-13
CRISTAL-13 (vuds_cosmos_5100994794, at , shown in Fig. 2, has the largest physical extent of the four CRISTAL galaxies studied in this paper. It is formed of a compact North East component, and a diffuse and clumpy tail to the West, which extends over 1.5 arcsec or 10 kpc. In the Western tail (WT), we identify at least 5 separate clumps in the higher spatial resolution F090W to F200W bands shown in Fig. 1. When the data is convolved to the resolution of the F444W band for the SED fitting analysis, these regions are less easily distinguished, however we see structure in the derived parameter maps over these scales (e.g. see the contours shown in the F090W-F150W map in Fig. 2). The clumps within the Western tail show correlated values of high sSFR, young age, and blue rest-frame UV slope, which are also correlated with the F090W–F150W colour as expected. We see several peaks in the derived H and H + [OIII] emission in the Western tail, further corroborating that these are regions of high sSFR with young ( Myr) ages.
The NE component and the Western tail are both very blue () and embedded within regions of redder rest-frame UV colour. We verified that this halo of redder colours was present in the images directly, and not an artefact of the fitting process. The NE component is compact and shows a colour gradient such that the galaxy is reddest between the NE component and the blue tail. This NE regions is characterised by a relatively complex age and structure, which is due to degeneracies between these properties in the fitting. The redder region of the NE component, which is also picked up in the Balmer (F150W–F277W) colour maps, corresponds to the centre of mass of the system, which is subtly offset from the peak in SFR. The peak in mass is however co-spatial with the peak in . Despite containing a quarter of the pixels, the wider NE component (that includes the SFR and stellar mass peaks) contains half of the total mass. CRISTAL-13 contains a significant ALMA detection of the dust continuum (8), which spatially coincides with the strongest inferred .